The force attraction between eva and mika is 2x10^2N. Mika has a mass of 80kg. What is the mass of Eva if they ate 200m apart?
1
Expert's answer
2020-12-03T07:00:35-0500

We can find the mass of Eva from the Newton's law of universal gravitation:


"F=G\dfrac{m_{Eva}m_{Mika}}{r^2},""m_{Eva}=\dfrac{Fr^2}{Gm_{Mika}},""m_{Eva}=\dfrac{2\cdot10^2\ N\cdot (200\ m)^2}{6.672⋅10^{−11}\ \dfrac{Nm^2}{kg^2}\cdot 80\ kg}=1.49\cdot10^{15}\ kg."

Answer:

"m_{Eva}=1.49\cdot10^{15}\ kg."